If f(x+1)=2x^(2)+x+a-1
f(3)=?
a=?
1
Expert's answer
2020-04-14T17:21:29-0400

"f(x+1)=2x^2+x+a-1"

"f(3)" will be when "x=2"

"f(3)=2\times2^2+2+a-1=9+a"

"a" can be any value, since we don't have more information